OpenStudy (anonymous):

graphing help please, MEDAL. A system of linear inequalities is shown below: x + 2y < -5 x - y > 0 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

someone?? @iGreen ?

OpenStudy (anonymous):

you'll want to start with putting it in y=mx+b form.

OpenStudy (anonymous):

i know, but i'm not sure how.

OpenStudy (anonymous):

Start with the first equation. We want to get rid of that 2 by cancelling it out. We divide the left side by 2 but what we do to one side we must do to the other. so we divide -5 by 2 to get -2.5 This makes our equation x+y<-2.5

OpenStudy (anonymous):

all we do with these equations now is move the x y<-x-2.5 -y>x

OpenStudy (anonymous):

we graph both of the intercepts.

OpenStudy (anonymous):

The first equations intercept is -2.5 and the second one is 0

OpenStudy (anonymous):

Since the slope is \[\left(\begin{matrix}1 \\ 1\end{matrix}\right)\] this means you go up one and over one for each one.

OpenStudy (anonymous):

x - y > 0 = -y > x ?

OpenStudy (anonymous):

it actually translates to -y>x-0 but -0 is non existent so can be removed.

OpenStudy (anonymous):

oh, okay ! thank you !

OpenStudy (anonymous):

No problem! :)